# GPTKB: Building Very Large Knowledge Bases from Language Models

### v2

@@ 4 Implementation
We chose GPT-4o-mini OpenAI (2024) for our experiments, because of its relatively good tradeoff between performance and cost, and its ability to process requests in batches. This means we can send multiple subjects from our queue at the same time, after startup, typically 5,000 entities per batch, and up to 100 batches in parallel. The model’s size is not publicly released, but has been estimated to be around 8B. [^3] See https://explodingtopics.com/blog/gpt-parameters, which in turn links to https://techcrunch.com/2024/07/18/openai-unveils-gpt-4o-mini-a-small-ai-model-powering-chatgpt, which references a statement by OpenAI that we could not find. But the underlying reasoning based on performance comparison with open models appears sensible.
@@
We start the process from Vannevar Bush, the visionary of interlinked knowledge management. We require a total of 2,200 batches to prompt GPT for 5.8M entities, obtaining in the answers triples for over 2.9M entities of them, which took us about 27 hours. The subject queue was still not empty at that point, but contained overwhelmingly duplicate or hallucinated subjects (see discussion below). Including trial runs, constructing GPTKB costed us a total of \$2,200 for OpenAI API calls.
@@
Our KB contains a total of 105M triples for 2.9M entities, organized into 2,133 relations and 367 classes. This gives an average of 36 triples per subject, with two distinct clusters, in particular, 651K entities with 10 triples and 86K with 50 triples, with most others near these values. Of all triples, more than 37M have an entity as object, and more than 67M have a literal as an object. The average length of entity labels is 24.5 characters. Example output for Vannevar Bush is shown in Table 2.
@@
We provide our KB both as a download (3.8 GB in TTL format), via a web browsing interface, and via a SPARQL query endpoint at http://gptkb.org. We provide our KB under the noncommercial CC BY-NC 4.0 license, not out of commercial self-interest, but because we want to avoid potential misuses (see Section 8). This license is compatible with the permissive terms of use of OpenAI. [^4] https://openai.com/policies/sharing-publication-policy
@@ 5 Analysis
5.1 GPTKB content
The most frequent classes generated for entities are shown in Fig. 5, where we find that patent and Person dominate the dataset. Several more classes are actually subclasses of Person, e.g., Noble, Duke, historical figure. Including these, the total number of humans is close to 600K. We also find a largely hallucinated class, US Coast Guard cutters, among the top-3.
@@ 5.1 GPTKB content
Figure 5: The 12 most frequent classes in GPTKB, which make up 32% of all entities.
Relation Object
instanceOf person
birthPlace Everett, Massachusetts
born March 11, 1890
affiliation Massachusetts Institute of Technology
United States government
influencedBy John von Neumann
knownFor Memex
awards National Medal of Science
Presidential Medal of Freedom
Table 2: Excerpt from 41 triples for subject Vannevar Bush in GPTKB. Entity objects for which further triples were found are highlighted in italics.
@@
The most frequent properties in the whole KB are patentCitation (3.15M) and instanceOf (2.96M). with similarly generic properties following at the KB level. At the class level, we find more specific properties, for instance, the most frequent properties in the class Person are instanceOf (251K), hasOccupation (126K), knownFor (119K), nationality (114K).
@@
We also analyze the geographic bias of our KB, using the listed nationalities as proxy. The results are shown in Fig. 6. We can observe a clear bias towards English-language nationalities, stronger than, e.g., in Wikidata Shaik et al. (2021), and a curious lack of others (e.g., Chinese only 2K), likely reflecting the (undisclosed) English-centric training corpus of GPT-4o-mini. Further interesting observations concern the LLM’s reluctance to express the gender property (only 30K triples, with female at 15K more present than male at 8K), likely reflecting the measures taken towards hiding sensitive attributes, and active debiasing.
[…]
Figure 6: Distribution of values for nationality.
5.2 Wikidata comparison
@@
We compare with Wikidata on several aspects: First, we compute the fraction of subjects that exist in Wikidata. For a random sample of 200 subjects from GPTKB, 24% have an entity with exactly matching label in Wikidata. A further 6.5% have a non-empty search result, i.e., an entity of paraphrased or similar label. The remaining 69.5% appear novel. Second, we exemplarily look at the 41 triples for Vannevar Bush, of which we find that more than 10 are not contained in Wikidata, e.g., his affiliation with the US government, his children count (incorrect by 1), or him inventing the concept of hypertext. [^5] An exact comparison is not straightforward, because relation names do not perfectly align, and objects in GPTKB are substantially more wordy. Third, we identify several properties not modelled at all in Wikidata, for instance, historicalSignificance (270K triples), hobbies (30K triples), or hasArtStyle (11K triples). This indicates that GPTKB potentially contains a significant amount of novel knowledge. A more comprehensive KB comparison in the style of Färber et al. (2018) is planned.
5.3 Precision and hallucinated content
Evaluating the precision of large-scale KBs is not straightforward, because they contain a significant amount of long-tail knowledge, for which finding evidence or counter-evidence is difficult. In line with previous work Suchanek et al. (2007), we evaluate precision within the context of web-retrievable information.
In particular, for a given triple, we retrieve the top-5 web-page snippets for the search term <subject object> from a search engine API, then use another LLM, namely Llama-3.1 llama31, to judge whether the given triple appears true/plausible/implausible/false, given the snippets. We adapt the prompt used by Adam and Kliegr (2024), extending it with the “false” level. From a random sample of 200 triples, we find that, based on the snippets, 22.5% appear true, 57.5% appear plausible, 1% appear implausible, 19% appear false. We also sample 200 triples from entities that are instances of Person, finding there that 21.5% of their triples appear true, 52.5% appear plausible, 0% appear implausible, 26% appear false. Extrapolating the former results, we can estimate that GPTKB contains approximately 24M triples that can be web-verified as true, 60M that appear plausible according to web context, 1M that appear implausible, and 19M that are contradicted by web context.
In terms of subjects, we observed a wide variance. While for some classes, e.g., Person or Company, most entities appeared real, others showed a strong runaway, e.g., patents, where most codes were made up (e.g., US_04,500,002_A1), or United States Coast Guard Cutter (the LLM continuously added varying incorrect type numbers to real names, e.g., USCGC_Biscayne_(WLB-201)). We therefore also evaluate the web verifiability of subjects alone. Similar to the approach used for triples, we employ a search engine to retrieve 5 snippets, now given only the entity’s name as query. We then evaluate three levels: (i) The snippets contain an entity of exactly that name, (ii) the snippets make it likely that the entity exists, perhaps under a slightly different label, (iii) the snippets give no support for the existence of that entity. From a random sample of 200 entities, we find that 41.5% fall into the first category, 43.5% into the second, 15% into the third, implying an extrapolated 1.2M verifiable, 1.3M plausible, and 0.4M questionable KB subjects. Among subjects classified as Person, 67% fall into the first category, 21.5% into the second, and 11.5% into the third. In the contrast, only 0.5% of subjects under United States Coast Guard Cutter fall into the first category, 13.5% into the second and 86% into the third.
About the accuracy of the taxonomy, we follow the per-edge evaluation scheme proposed by Bordea et al. (2016), finding that 64% of all subclass-superclass edges are considered correct, using another LLM as judge. We also evaluate whether a superclass is most appropriate, by offering all siblings as alternatives. Here, for 70% of all subclass-superclass edges, the superclass is considered the best alternative. Note that the structure of our taxonomy still leaves room for improvement in terms of long-range dependencies and distributions, that are not easy to quantify or address locally.
@@ 6.1 Lessons for LLM epistemiology
LLMs and the way they store, retrieve, and generate knowledge are under intense scrutiny from computational, as well as cognitive and philosophical perspectives Nosta (2024). We have shown that one can iteratively materialize internal knowledge of an LLM at an unprecedented scale, but the notion and definition of LLM knowledge itself is controversial Fierro et al. (2024). We adopt here the term knowledge base because it is common in the field, but would consider the term belief base equally appropriate. In terms of the definitions of knowledge that are introduced by Fierro et al. (2024), our output falls into the minimalistic sui-generis (g-knowledge) category, as our KB may contain false as well as inconsistent triples.
We can now, in principle, relate the number of parameters of an LLM to the amount of entity-centric triple knowledge it holds: As discussed in Section 4, GPT-4o-mini likely has in the order of 8B parameters. Given the 105M triples that we obtained, that makes 76 parameters per triple. [^6] From Allen-Zhu and Li (2024), a theoretical lower bound of 4 parameters/triple can be deduced, based on Remark 4.4, and their synthetic dataset containing 6 triples per entity. However, the generalizability of results on that synthetic dataset is unclear. We emphasize that this number captures only encyclopedic knowledge, yet that LLMs also possess other knowledge, e.g., linguistic knowledge, procedural knowledge, etc.
As already observed by Petroni et al. (2019), prompt formulations heavily influence LLM responses, and hence may bias the resulting KB. GPTKB presents one way of materializing LLM knowledge into structured format, but it is important to observe that LLM knowledge is inherently probabilistic, while our output (and most web-scale KBs) are binary. Any process translating from probabilistic to binary implicitly encodes a choice between precision and recall, and we do so by the triple counts given as guidance in the elicitation prompt. In our view, such a biasing is inherently unavoidable, and also not solved by instruction-free sample-only (few-shot) prompts as proposed by Wu et al. (2024b), as also examples bias the responses, both in terms of quantity, topics, and vocabulary.
@@ 6.2 Lessons for KB construction
Our exploration surfaced a potpourri of other challenges, some of them known to KB construction research since years Weikum et al. (2021), others requiring novel adaptations in the light of LLMs. These concern, in particular, (1) NER for short labels without context, (2) entity deduplication, (3) entity canonicalization, (4) literal typing and canonicalization, (5) relation clustering, (6) relation organization in terms of subrelations, (7) class clustering, (8) taxonomy construction, and (9) triple verification. While we explored simple techniques for 1, 2, 5, 7 and 8, each has room upward, and the other tasks were not treated so far.
@@
The GPTKB approach deviates from traditional Wikimedia- and data-integration focused KB construction, and although its precision still needs improvement, it stands out with its potential for cost efficiency. In a back-of-the-envelope calculation, Paulheim (2018) estimated the cost per triple for existing manual KB construction projects at \$2-6, and for existing automated KB construction projects at \$0.01-0.15. In contrast, in our prototypical execution, the cost is just \$0.00009 per correct triple (\$2.2K/24M triples), i.e., over 100x less than with previous methods.
@@ 6.3 Other LLMs
We performed a parallel exploration using Llama 3.1 70B on local HPC hardware, but found this not competitive in terms of runtime. We also envisioned a run using the strongest publicly available LLM, GPT-4o, however, by its 15x higher API cost, and its estimated 25x larger parameter set, assuming knowledge is roughly proportional to parameters, at about \$825K, this is beyond our budget.
@@ 7 Conclusion
We proposed to build large general-domain KBs entirely from LLMs. We showed that large-scale KB construction from LLMs is feasible, while highlighting specific challenges arising around precision, and KB construction tasks like entity recognition, entity and property canonicalization, and taxonomy construction. As a prototype, we used GPT-4o-mini to construct GPTKB, which contains 105 million assertions for over 2.9 million entities, at a cost 100x cheaper than existing projects.
Our work is a landmark for two fields: For NLP, for the first time, it provides constructive insights into what LLMs know (or believe). For the Semantic Web, it shows novel ways forward for the long stale topic of open-domain KB construction. GPTKB is accessible at http://gptkb.org.
@@ 8 Limitations
GPTKB presents a proof of concept that so far does not satisfy the precision requirements of most downstream use cases, nor is its precision even accurately established, because it contains many long-tail triples without web sources that enable confirming or refuting their truth. The knowledge in our KB can intrinsically not be sourced back to any specific document and contributor. Also, as we do not prescribe any standardized vocabulary, and only apply lightweight consolidation, our KB does not yet possess the consistency of existing projects. If you consider using a KB in production, for the time being, consider an alternative like Wikidata.
As there are different ways of how an LLM can be prompted, GPTKB, based at its core on one knowledge elicitation prompt, can only provide one constructive materialization of LLM knowledge, but others with different content and recall or precision exist (see Section 6.1 on “LLM epistemiology”).

### v3

@@ 4 Implementation
We chose GPT-4o-mini OpenAI (2024) for our experiments, because of (i) its ability to process requests in batches, (ii) its structured output feature, (iii) its good tradeoff between performance and cost. The batch feature is particularly important, enabling us to send, after startup, up to 100 batches of 10,000 entities in parallel. The model’s size is not publicly released, but has been estimated to be around 8B Howarth (2024); Zeff (2024).
@@
We start the process from a single entity, Vannevar Bush, in honor of his vision of interlinked knowledge Bush (1945). Note that this choice is arbitrary and overall inconsequential. As general-world knowledge graphs are densely connected Hogan et al. (2022), any seed entity connected to popular entities serves as well. For instance, we reach MIT in one hop, Alan Turing in 2 hops, Kyoto in 3 hops.
We require a total of 2,200 batches to prompt the LLM up to BFS-depth 10, which corresponded to 5.8M prompted entities, of which for 2.9M a non-empty answer was returned. The whole process took 27 hours, and including trial runs, constructing GPTKB cost us a total of \$3,500 for API calls.
@@
Our KB contains a total of 101M triples for 2.9M entities, organized into 1,804 relations and 473 classes. This gives an average of 36 triples per subject, with two distinct clusters, in particular, 651K entities with 10 triples and 86K with 50 triples, with most others near these values. Of all triples, more than 37M have an entity as object, and more than 67M have a literal as an object. The average length of entity labels is 24.5 characters. Example output for Vannevar Bush is shown in Figure 2. Statistics are shown in Table 3.
Entities 2.9M
Triples 101M
Relations 1,804 (788K before canonicalization)
Classes 473 (103K before canonicalization)
Triple objects 37M entities, 64M literals
Avg. triples/entity 36
Avg. label length 24.5 characters
Avg. outlinks 4
Subject-precision* 74% Verifiable, 9% Plausible,
17% Unverifiable
Subjects in Wikidata* 37% in WD, 63% not in WD
Triple-precision* 31% True, 61% Plausible,
1% Implausible, 7% False
* as a weighted average across layers.
Table 3: Statistics of GPTKB.
@@
We provide our KB both as a download (3.8 GB in TTL format), via a web browsing interface, and via a SPARQL query endpoint at https://gptkb.org. We provide our KB under the permissive CC BY-NC 4.0 license, though we caution against unreflected downstream usage (see Section 8). This license is compatible with the permissive terms of use of OpenAI OpenAI (2022).
@@ 5 Analysis
5.1 Precision
Evaluating the precision of large-scale KBs is not straightforward, because they contain a significant amount of long-tail knowledge, for which finding evidence or counter-evidence is difficult. In line with Suchanek et al. (2007), we evaluate precision within the context of web-retrievable information. Figure 3 summarizes our findings.
For verifying entities, we retrieved 5 search results from a web search API for the entity label, which are then used to decide whether the entity’s existence could be verified, appears plausible, or could not be verified. Deciding on these labels based on textual context represents a task of textual entailment, a.k.a. natural language inference (NLI), a task generally considered solved for LLMs. [^2] On LLM performance on textual entailment, the popular SNLI Bowman et al. (2015) is comparable, which was considered solved in 2020 by BERT architectures Zhang et al. (2020) at 92% accuracy, which is the ceiling, given dataset noise. This is also why the literature mentions no numbers for newer models like Llama on this task anymore. We therefore use another LLM, Llama-3.1-70B-Instruct Dubey et al. (2024), for this task. The LLM hereby only judges textual entailment, it does not independently judge truth from its parameters. Based on 1,000 samples, average entity verifiability is 74%, and as shown in Figure 3, we observe a continuous drop of verifiability over the layers, stabilizing at layer 10 at about 70%. The plot also includes Wikidata-existence, which paints a similar trend at lower levels. We generally observe a wide variance across classes, finding, for instance, that persons consistently have a higher verifiability than fictional character, where many codes were made up (e.g., Officer K.I.T.T. XV and Officer K.I.T.T. XVI) or varying incorrect type numbers are added to real names continuously by the LLM (e.g., Agent 71 and Agent 72).
111122223333444455556666777788889999101010100020202020404040406060606080808080100100100100BFS-LayerPercentage (%)Entities web-verifiableEntities in WikidataTriples web-verifiableTriples web-plausible
Figure 3: Accuracy of GPTKB across layers.
For a given triple, we retrieve the top-5 web-page snippets for the search term <subject> <object> from a search engine API, then again perform LLM-based textual entailment to decide whether the given triple is entailed (entailed/plausible/implausible/false), given the snippets. We specifically adapt the prompt used by Adam and Kliegr (2024), extending it with the “false” level. Based on 1,000 samples, we find that on average, 31% of triples are entailed, 61% are plausible, 1% is implausible, and 7% are false.
In terms of the accuracy of the taxonomy, we follow the per-edge evaluation scheme proposed by Bordea et al. (2016), finding that 64% of all subclass-superclass edges are considered correct, using the Llama model as judge. We also evaluate whether a superclass is most appropriate, by offering all siblings as alternatives. Here, for 70% of all subclass-superclass edges, the superclass is considered the best alternative. Note that the structure of our taxonomy still leaves room for improvement in terms of long-range dependencies and distributions, that are not easy to quantify or address locally.
5.2 GPTKB content and comparison
Classes
The most frequent classes generated for entities are shown in Figure 4, where we find that Person dominates the dataset. Several more classes are actually merged into Person, e.g., human, hominid. Including these, the total number of Person is more than 287K. Further details by layer are in Appendix B.
@@ 5.2 GPTKB content and comparison
Figure 4: The 10 most frequent classes in GPTKB, which make up 25% of all entities.
Properties
@@
The most frequent properties in the whole KB are instanceOf (3.1M) and features (2M) with similarly generic properties following at the entity level. At the class level, we find more specific properties, for instance, the most frequent properties in the class Person are instanceOf (309K), occupation (198K), knownFor (179K), exhibition (178K).
Content bias
@@
We also analyze the geographic bias of our KB, using the listed nationalities as proxy. We observe a clear bias towards English-language nationalities: the top-2 are American (119K) and British (35K), while the next 3 are French (18K), German (14K), and Japanese (11K). This bias in GPTKB is stronger than, e.g., in Wikidata Shaik et al. (2021), and a curious lack of others (e.g., Chinese only 3K, compared to 8K for Russian, or 7K for Indian/Spanish), likely reflecting the (undisclosed) English-centric training corpus of GPT-4o-mini.
An interesting point is also gender bias: The gender property of GPTKB contains 15K female versus 8K male values. For first names, male ones are still more than female (47% vs. 37% based on the gender-guesser 0.4.0 Python library), but this is still a much lower bias than in other KBs, likely reflecting the effects of gender debiasing.
Wikidata overlap and novelty
@@
We compare with Wikidata on several aspects: First, we compute the fraction of subjects that exist in Wikidata. From a random sample of 2,000 subjects from GPTKB, 24% have an entity with exactly matching label in Wikidata. A further 6.5% have a non-empty search result, i.e., an entity of paraphrased or similar label. The remaining 69.5% appear novel. Second, we exemplarily look at the 41 triples for Vannevar Bush, of which we find that more than 10 are not contained in Wikidata, e.g., his affiliation with the US government, his children count (incorrect by 1), or him inventing the concept of hypertext. [^3] An exact comparison is not straightforward, because relation names do not perfectly align, and objects in GPTKB are substantially more wordy. Third, we identify several properties not modelled at all in Wikidata, for instance, historicalSignificance (396K triples), artStyle (174K triples), or hobbies (44K triples). This indicates that GPTKB potentially contains a significant amount of novel knowledge. A more comprehensive KB comparison in the style of Färber et al. (2018) is planned.
@@ 6.1 Lessons for LLM epistemiology
Terminological observations
The notion and definition of LLM knowledge itself is controversial Fierro et al. (2024). We adopt here the term knowledge base because it is common in the field, but would consider the term belief base equally appropriate. In terms of the definitions of knowledge that are introduced by Fierro et al. (2024), we find that our output falls into the minimalistic sui-generis (g-knowledge) category, as our KB contains false as well as inconsistent triples. Beyond the reported errors, e.g., in many cases, spouse relations are not symmetrically represented.
Storage capacity
Scaling laws and storage capacity are intensively investigated Allen-Zhu and Li (2024), but usually on synthetic data. Our experiments provide a lower bound on real data: As discussed in Section 4, GPT-4o-mini likely has in the order of 8B parameters. Given the 101M triples that we obtained, that makes 79 parameters per triple. [^4] From Allen-Zhu and Li (2024), a theoretical lower bound of 4 parameters/triple can be deduced, based on Remark 4.4, and their synthetic dataset containing 6 triples per entity. However, the generalizability of results on that synthetic dataset is unclear. We emphasize that this number captures only encyclopedic knowledge, yet that LLMs also possess other knowledge, e.g., linguistic knowledge, procedural knowledge.
Complementarity
We observe a high complementarity to existing resources, with 63% of generated entities not being present in Wikidata. Particularly, from Figure 3, GPTKB’s entities remain web-verifiable even in deeper layers, about 70% of the entities are web-verifiable in layer 10, even though the percentage of entities found in Wikidata drops to a little over 20%. Apart from named-entities, some particularly noteworthy complementary slices of LLM knowledge concern digital media artifacts, art periods, and people’s hobbies.
Accuracy
While on the entity level, most entities appear to truly exist (74% verifiable, 9% plausible), on the triple level, results are lower (31% verifiable, 61% plausible). This indicates that the LLM generally has a grasp of entities, but has more difficulties in correctly modeling their relationships. BFS-depth and quality are negatively correlated, but there is no complete quality collapse within the first 10 levels.
Enabled analyses
Our materialized resource enables a range of analyses about factual LLM knowledge, which for space reasons we only skim here: 1) Accuracy: See Sec. 5.1 and paragraph above. 2) LLM bias: See Sec. 5.2. 3) Timeliness: Exemplarily, we collect the most frequently mentioned years, and observe that there is sharp drop after 2023, which matches the knowledge cutoff of the LLM (see plot in Appendix D). 4) Subject-wise consistency: Our KB still contains many duplicates, e.g., The Elbe River, River Elbe, Elbe River, river Elbe, Elbe. Studying their triples gives insights into consistency, e.g., we observe a significant semantic overlap, but also frequent different wordings (e.g., wildlife - various fish species / fish species), and minor factual deviations (e.g., length 3x 1094km, 2x 1091km). 5) Structural consistency: An interesting observation concerns the difficulty that LLMs have with inverse relations Berglund et al. (2024). We can observe this too, for instance, out of 318K spouse triples, only 8K are symmetric, and out of 61K parentCompany triples, only 6K are mirrored in subsidiary triples.
@@ 6.2 Lessons for KB construction
Emerging entities
Some applications are especially interested in newly emerging entities, and these are a long-standing challenge for traditional KBC Hoffart et al. (2016). Web-scraped KBs like Yago and DBpedia could in principle re-run their scrapers periodically, while text-extracted KBs like Nell and ReVerb would require re-runs on new web crawls. Our approach could proceed analogous, re-running the materialization on a newer version of the utilized LLM. If one were to know which entities are affected by updates, one could also perform retrieval-augmented generation, however, knowing where updates occured essentially requires an oracle, and none of the existing KBs successfully employed selective updating.
@@
Our exploration surfaced a potpourri of other challenges, some of them known to KB construction research since years Weikum et al. (2021), others requiring novel adaptations in the light of LLMs. These concern (1) NER for short labels without context, (2) entity deduplication, (3) entity canonicalization, (4) literal typing and canonicalization, (5) relation clustering, (6) relation organization in terms of subrelations, (7) class clustering, (8) taxonomy construction, and (9) triple verification. While we explored simple techniques for 1, 2, 5, 7, 8, each has room upward, and other tasks were not treated so far.
@@
The GPTKB approach deviates from traditional Wikimedia- and data-integration focused KB construction, and although its precision still needs improvement, it stands out with its potential for cost efficiency. In a back-of-the-envelope calculation, Paulheim (2018) estimated the cost per triple for existing manual KB construction projects at \$2-6, and for existing automated KB construction projects at \$0.01-0.15. In contrast, in our prototypical execution, the cost is just \$0.0001 per correct triple (\$3.5K/33M), i.e., $>$100x less than with previous methods.
@@ 6.3 Other LLMs
We performed a parallel exploration using Llama-3.1-70B-Instruct on local HPC hardware, and while accuracy is higher (69% of triples verified true in a test run), this did not scale. We also envisioned a run using the strongest publicly available LLM, GPT-4o (80% triples verified true), however, by its 15x higher API cost, and its estimated 25x larger parameter set, assuming knowledge is roughly proportional to parameters, at about \$825K, this is beyond our budget. See Appendix E.
@@ 7 Conclusion
Our work is a landmark for two fields: For NLP, for the first time, it provides comprehensive insights into what LLMs know (or believe). For the Semantic Web, it shows novel ways forward for the long stale topic of open-domain KB construction. GPTKB is accessible at https://gptkb.org.
@@ 8 Limitations
As already observed by Petroni et al. (2019), prompt formulations heavily influence LLM responses, and hence may bias the resulting KB. GPTKB presents one way of materializing LLM knowledge into structured format, but different prompts could give different output. It is also important to observe that LLMs operate inherently probabilistically, while our output (and most web-scale KBs) are binary. Any process translating from probabilistic to binary implicitly encodes a choice between precision and recall, and we do so by the triple counts given as guidance in the elicitation prompt. In our view, such a biasing is hard to avoid, and also not solved by instruction-free sample-only (few-shot) prompts as proposed by Wu et al. (2024), as also examples bias the responses, both in terms of quantity, topics, and vocabulary.
Our exemplary run presents a proof of concept for materializing LLM knowledge. While this knowledge has slices complementary to existing KB knowledge, so far our experiments on verifiability does not satisfy the precision requirements of most downstream use cases, nor is its precision even accurately established, because it contains many long-tail triples without web sources that enable confirming or refuting their truth. The knowledge in our KB can intrinsically not be sourced back to any specific document and contributor. Also, as we do not prescribe any standardized vocabulary, and only apply lightweight consolidation, our KB does not yet possess the consistency of existing projects. If you consider using a KB in production, for the time being, consider an alternative like Wikidata.
@@
Acknowledgment
We thank Markus Krötzsch and Gerhard Weikum for comments on this work.